How long (seconds) will it take the dime to fall 4.9 meters?
Answer by mananth(16949)

S= ut +1/2 at^2
S=4.9
a=9.8 m/s^2
4.9=0+1/2*9.8*t^2
u=0 since it is dropped
4.9/4.9=t^2
1=t^2
t=1 second